OpenStudy (anonymous):

n a 45°- 45°- 90° right triangle, the hypotenuse is 5 cm. How long are the legs?

jimthompson5910 (jim_thompson5910):

Let x = length of each leg (they're both equal to each other) So a^2 + b^2 = c^2 x^2 + x^2 = 5^2 2x^2 = 25 I'll let you finish

jimthompson5910 (jim_thompson5910):

I'm using the pythagorean theorem in my first step if you're curious

OpenStudy (anonymous):

So ill juss divide each side by 4 because you times 2 times itsself ?

jimthompson5910 (jim_thompson5910):

no you divide each side by 2

jimthompson5910 (jim_thompson5910):

the squared only applies to the x

OpenStudy (anonymous):

ohh ok..

OpenStudy (anonymous):

Each leg is the same length,\[5 \text{ Sin}[45{}^{\circ}]=\frac{5}{\sqrt{2}}=3.53553 \]

OpenStudy (anonymous):

Yeah , thats what i got ,, thank you .

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

review this and you will gain the proper knowledge in your future mathematical endeavor http://www.purplemath.com/modules/specang.htm